# Contrastive Diffusion Guidance for Spatial Inverse Problems

### v1

@@ 4 Experiments and Evaluation
Baselines. We evaluate the performance of CoGuide against 6 competitive baselines: $\blacksquare$ DPS+X: DPS Chung et al. (2022) based inverse solver with 3 differentiable path-planners $\mathcal{A}$ by instantiating $\textbf{X}\!\in\!\{\text{NeuralA*},\text{TransPath},\text{DiPPeR}\}$ Yonetani et al. (2021), Kirilenko et al. (2023), Liu et al. (2024). More details about the planners are provided in the Appendix D. Briefly, these path-planners use a CNN-based shortest-path module, a Transformer based path-probability encoder, and a diffusion-based planner, respectively. $\blacksquare$ DiffPIR Zhu et al. (2023): A plug-and-play image restoration solver that uses a diffusion model as a denoiser, instead of training a Gaussian denoiser. We follow the default configuration in https://github.com/yuanzhi-zhu/DiffPIR. Since we do not have a closed form data proximal estimator, we use an ADAM optimizer to solve the proximal with $10$ optimization steps and use a $0.01$ learning rate. $\blacksquare$ DMPlug Wang et al. (2024): A recent inverse solver that optimizes the noise seed such that after a DDIM sampler, the resulting image satisfies the measurement constraint. We follow the default configurations in https://github.com/sun-umn/DMPlug except we use 100 optimization steps. $\blacksquare$ CFG: Performs Classifier-free Guidance Dhariwal & Nichol (2021) that combines the unconditional and trajectory-conditioned scores with a guidance scale that biases sampling toward trajectory-consistent floorplans.
Diffusion and Contrastive Model. Our base diffusion model follows the implementation used in the DPS Chung et al. (2022). We adapted the contrastive model used in Khosla et al. (2020) based on the specific needs of this project. We discuss architecture and hyperparameter settings in Appendix C.
@@ 4.1 Results
$\blacksquare$ Quantitative. Table 1 reports F1/IoU (mean $\pm$ std) for three regimes of trajectory density: sparse, moderate and dense. In the sparse regime, CoGuide attains the best performance, exceeding all baselines, including CFG. Similarly in the moderate regime, CoGuide again leads surpassing CFG and the DPS variants. However, in the dense case, CFG is the strongest, with CoGuide showing comparable performance. Overall, CoGuide consistently outperforms DPS+X, DiffPIR, and DMPlug.
@@ 4.2 Ablations
As discussed in section 3.3, we incorporate the Adam update during the sampling process and compare against the standard Gradient descent-based update. Results reported in Table 4.2 shows that employing Adam consistently outperforms SGD across both DDPM and DDIM.
$\blacksquare$ Measurement Noise. In real-world settings, the localization sensors that provide the input trajectories to CoGuide may be noisy. To model this, we inject Gaussian noise into the trajectory generation process. We increase the noise standard deviation and compare our performance across various noise levels and trajectory densities. CoGuideś performance understandably degrades with increasing noise levels as shown in Fig. 4.2. The degradation is however graceful, and denser trajectory upholds better performance.
@@ 5 Related Work
Other modalities have also been explored beyond vision, including acoustics Zhou et al. (2017), magnetics Luo et al. (2017), RF Peng et al. (2018), radar Hickman & Krolik (2009). While effective in specific settings (e.g., BatMapper Zhou et al. (2017)), many require specialized hardware or calibrated infrastructure, whereas IMU-based trajectories are easy to obtain. Walk2map Mura et al. (2021) is the most relevant to our work, but it is designed for single-room layouts and does not have any generative capabilities.
6 Follow-On Work and Conclusion
$\blacksquare$ Spatial Inverse Problems: This paper focused on a specific floorplan estimation problem, however, the notion of contrastive guidance should lend itself to a broader family of non-differentiable $\mathcal{A}$ operators. We intend to investigate what family of operators can benefit, and conversely, how can the contrastive guidance be improved to broaden that family of operators. Along these lines, newer applications are also of interest. For instance, can city maps be synthesized based on GPS trajectories of vehicles? Can discrete molecular structure be synthesized from measured properties of molecules? Can Internet topologies be derived based on streaming packet analytics? $\blacksquare$ Towards Realism: Even in our specific floorplan application, there is room for improvement. The standard floorplan dataset (HouseExpo) does not include furniture. Incorporating furniture into the environment is a practical extension, if such a dataset is available. $\blacksquare$ Blind Inverse Problems: That the contrastive loss is almost agnostic to the $\mathcal{A}(.)$ operator may enable approaches to blind inverse problems, if we can generate measurement $\mathbf{y}$ from known $\mathbf{x}$’s. If partial information is available about $\mathcal{A}(.)$, could that be adequate to design a good embedding space? We believe CoGuide could initiate conversation along all these branches, inviting a range of follow-on research and exploration.
@@ Appendix A Likelihood under InfoNCE objective
Matching $q$ and $p$ requires that the numerators differ only by a multiplicative constant (since softmax is invariant to shifts). Thus at the InfoNCE optimum,
@@ Appendix B Additional Qualitative Results
$\blacksquare$ Additional Results. We show results from evaluation on additional floorplans in Fig. 5 and 6. We observe that CoGuide significantly outperforms all inverse-solver baselines while being better or comparable to CFG.
@@ C.1 Diffusion Model
C.2 Supervised Contrastive Learning
@@
C.3 Loss function descriptions
@@
Appendix D Path-planning algorithms

### v2

@@ 4 Experiments and Evaluation
Blind Inverse Problems. CoGuide extends naturally to blind inverse problems. We demonstrate this in Appendix D through restoration of historical audio recordings containing fully unknown degradations.
@@
Baselines. We evaluate the performance of CoGuide against 6 competitive baselines: $\blacksquare$ DPS+X: DPS Chung et al. (2022) based inverse solver with 3 differentiable path-planners $\mathcal{A}$ by instantiating $\textbf{X}\!\in\!\{\text{NeuralA*},\text{TransPath},\text{DiPPeR}\}$ Yonetani et al. (2021), Kirilenko et al. (2023), Liu et al. (2024). More details about the planners are provided in the Appendix E. Briefly, these path-planners use a CNN-based shortest-path module, a Transformer based path-probability encoder, and a diffusion-based planner, respectively. $\blacksquare$ DiffPIR Zhu et al. (2023): A plug-and-play image restoration solver that uses a diffusion model as a denoiser, instead of training a Gaussian denoiser. We follow the default configuration given in their publicly available codebase. Since we do not have a closed form data proximal estimator, we use an ADAM optimizer to solve the proximal with $10$ optimization steps and use a $0.01$ learning rate. $\blacksquare$ DMPlug Wang et al. (2024): A recent inverse solver that optimizes the noise seed such that after a DDIM sampler, the resulting image satisfies the measurement constraint. We follow the default configurations in their publicly available codebase except we use 100 optimization steps. $\blacksquare$ CFG: Performs Classifier-free Guidance Ho & Salimans (2022) that combines the unconditional and trajectory-conditioned scores with a guidance scale that biases sampling toward trajectory-consistent floorplans.
Real-world Evaluation. To evaluate real-world performance, we collected trajectories in two student apartments using the Qorvo DWM3000 ultrawideband (UWB) module Qorvo (2021). The UWB system consists of multiple anchors statically placed at known 3D coordinates in the apartment. The user carries a beacon that periodically pings each anchor, allowing us to estimate the time-of-flight (ToF) to that anchor. Using the ToF to line-of-sight anchors and the known speed of light, we apply trilateration to estimate the user’s 3D position. We then project the 3D position onto the 2D floorplan to obtain walking trajectories. The UWB setup provides locations with a standard deviation of 10 cm. It is worth noting that the trajectories obtained via this method are sparser and distributed somewhat differently compared to our A*-generated dataset.
Diffusion and Contrastive Model. Our base diffusion model follows the implementation used in the DPS Chung et al. (2022). We adapted the contrastive model used in Khosla et al. (2020) based on the specific needs of this project. Classifier-free guidance (CFG) from the diffusers library is implemented by training a single UNet denoiser that can operate in both conditional and unconditional modes, and then combining its two outputs at inference with a guidance scale. The underlying idea and architecture come from the original classifier-free guidance paper Ho & Salimans (2022). We discuss architecture and hyperparameter settings for all our models in Appendix C.
@@ 4.1 Results
$\blacksquare$ Quantitative. Table 1 reports F1/IoU (mean $\pm$ std) for three regimes of trajectory density: sparse, moderate and dense. In the sparse regime, CoGuide attains the best performance, exceeding all baselines, including CFG. Similarly in the moderate regime, CoGuide again leads surpassing CFG and the DPS variants. However, in the dense case, CFG is the strongest, with CoGuide showing comparable performance. Overall, CFG+CoGuide consistently outperforms CFG, DPS+X, DiffPIR, and DMPlug.
@@
CFG+CoGuide $\mathbf{0.93\pm 0.04}$ $\mathbf{0.87\pm 0.07}$ $\mathbf{0.97\pm 0.03}$ $\mathbf{0.93\pm 0.04}$ $\mathbf{0.99\pm 0.01}$ $\mathbf{0.97\pm 0.02}$
@@ 4.2 Ablations
As discussed in section 3.3, we incorporate the Adam update during the sampling process and compare against the standard Gradient descent-based update. Results reported in Table 4.2 show that employing Adam consistently outperforms SGD across both DDPM and DDIM.
$\blacksquare$ Measurement Noise. In real-world settings, the localization sensors that provide the input trajectories to CoGuide may be noisy. To model this, we inject Gaussian noise into the trajectory generation process. We increase the noise standard deviation and compare our performance across various noise levels and trajectory densities. CoGuide’s performance understandably degrades with increasing noise levels as shown in Fig. 4.2. The degradation is, however, graceful, and denser trajectory upholds better performance.
@@
4.3 Real-world Floorplans
We compare CoGuide and CFG on real-world floorplans using UWB measured trajectories. Although both methods make errors, CoGuide substantially infers more wall segments and room structure, whereas CFG often misses or misplaces major walls (Fig. 5). This gap is expected since CFG is explicitly trained to model the posterior density under the synthetic A*-generated trajectory distribution, and therefore struggles to generalize when faced with the sparser, differently distributed real-world trajectories.
Additional inference results on more floorplans and uncertainty analysis are provided in Appendix B. Finally, in Appendix D we also demonstrate generalization of CoGuide beyond spatial inverse problems by tackling a blind audio enhancement application. Audio demos are provided on our website.
@@ 5 Related Work
Other modalities have also been explored beyond vision, including acoustics Zhou et al. (2017), magnetics Luo et al. (2017), RF Peng et al. (2018); Amballa et al. (2025), radar Hickman & Krolik (2009). While effective in specific settings (e.g., BatMapper Zhou et al. (2017)), many require specialized hardware or calibrated infrastructure, whereas IMU-based trajectories are easy to obtain. Walk2map Mura et al. (2021) is the most relevant to our work, but it is designed only for single-room floorplan layouts and does not have any underlying generative capabilities.
Figure 5: Real-world results comparing CFG and CoGuide on two apartment layouts. CoGuide reconstructed floorplans are qualitatively superior to those of CFG, which shows spurious structures.
6 Conclusion and Follow-On Work
$\blacksquare$ Spatial Inverse Problems: This paper focused on a specific floorplan estimation problem, however, the notion of contrastive guidance should lend itself to a broader family of non-differentiable $\mathcal{A}$ operators. We intend to investigate what family of operators can benefit, and conversely, how can the contrastive guidance be improved to broaden that family of operators. Along these lines, newer applications are also of interest. For instance, can city maps be synthesized based on GPS trajectories of vehicles? Can discrete molecular structure be synthesized from measured properties of molecules? Can Internet topologies be derived based on streaming packet analytics? $\blacksquare$ Towards Realism: Even in our specific floorplan application, there is room for improvement. The standard floorplan dataset (HouseExpo) does not include furniture. Incorporating furniture into the environment is a practical extension, if such a dataset is available. $\blacksquare$ Blind Inverse Problems: This paper showed early results on solving blind inverse problems in the audio domain. We believe this offers promise that contrastive guidance can enable a line of attack for complicated blind inverse problems, as long as we can generate measurement $\mathbf{y}$ from known $\mathbf{x}$’s. If partial information is available about $\mathcal{A}(.)$, could that be adequate to design a good embedding space? We believe CoGuide could initiate conversation along all these branches, inviting a range of follow-on research and exploration.
Acknowledgments
We are grateful to Sahil Bhandary Karnoor for helping us collect real-world data. We thank the anonymous reviewers for their valuable feedback. This work was partially supported by NSF #2008338, #1909568, #2148583, and #MRI-2018966. This work used DELTA at NCSA through allocation CIS230230 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296
@@ 7 Reproducibility Statement
    Generalizing CoGuide to Blind Inverse Problems
    1. (a)
      Blind Audio Restoration
    2. (b)
      Method
    3. (c)
      Contrastive Model Training and Implementation
    4. (d)
      Results
  5. 5.
@@ Appendix A Likelihood under InfoNCE objective
Matching $q$ and $p$ requires that the numerators differ only by a multiplicative constant (since softmax is invariant to shifts). Thus, at the InfoNCE optimum,
@@ Appendix B Additional Qualitative Results
$\blacksquare$ Additional Results. We show results from evaluation on additional floorplans in Fig. 6 and 7. We observe that CoGuide significantly outperforms all inverse-solver baselines while being better or comparable to CFG.
@@ C.1 Diffusion Model
C.2 Classifier-free Guidance
Architecture. Our diffusion prior uses a U-Net backbone with four resolution stages. The network operates on $64\times 64$ inputs and increases its channel capacity across levels $(128,256,384,512)$. Each resolution stage contains a single residual block, and a standard mid-block is placed between the downsampling and upsampling paths. The upsampling hierarchy mirrors the downsampling path to preserve multi-scale information. We use group normalization with 32 groups throughout the network, and time embeddings are injected via scale–shift modulation in each residual block so that the denoising behavior is explicitly conditioned on the diffusion timestep.
Training. We train the model for up to $50$ epochs on $64\times 64$ images with a batch size of $128$. The learning rate is set to $10^{-4}$ with $500$ warmup steps and no gradient accumulation. Classifier-free guidance Ho & Salimans (2022) is enabled by randomly dropping the conditioning with probability $0.1$ during training, so the network learns both conditional and unconditional denoising behavior in a single model. Training is stopped at $50$ epochs because the validation curve plateaus.
Inference and guidance tuning. At inference time, we run a standard diffusion sampler starting from Gaussian noise and iteratively denoising toward a clean floorplan. Given unconditional and conditional denoiser outputs, $s_{\theta}(\mathbf{x}_{t},t,\varnothing)$ and $s_{\theta}(\mathbf{x}_{t},t,\mathbf{c})$, we form the classifier-free guided prediction
$s_{\theta,\text{CFG}}(\mathbf{x}_{t},t,\mathbf{c})\;=\;(1+\gamma)\,s_{\theta}(\mathbf{x}_{t},t,\mathbf{c})\;-\;\gamma\,s_{\theta}(\mathbf{x}_{t},t,\varnothing),$ (18)
where $\gamma$ is the guidance scale. We tune $\gamma$ by sweeping over several values and selecting the setting $\gamma=1.5$ that achieves the best validation performance (Fig. 10).
Figure 10: Tuning the guidance for inference in CFG
C.3 Supervised Contrastive Learning
@@
C.4 Loss function descriptions
@@
Appendix D Generalizing CoGuide to Blind Inverse Problems
CoGuide’s contrastive diffusion-guidance enables us to solve blind inverse problems. We apply it to a blind audio restoration task on historical piano recordings that are degraded by unknown operators $\mathcal{A}(\cdot)$ (including spectral coloration, additive noise and clicks, nonlinear effects, etc.). This setting is fully blind because we neither assume an analytical form for $\mathcal{A}(\cdot)$ nor estimate it at inference time.
D.1 Formulation.
Like in CoGuide, we formulate this problem with a forward model. Here an unknown clean audio waveform $\mathbf{x}$ is observed through an unknown degradation operator $\mathcal{A}$ plus noise $\mathbf{w}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$:
$\mathbf{y}$ $=\mathcal{A}(\mathbf{x})+\mathbf{w}.$ (19)
Neither $\mathbf{x}$ nor $\mathcal{A}$ is known at test time. The goal is to recover a plausible $\hat{\mathbf{x}}$ that both explains the observation $\mathbf{y}$ and lies on the manifold of realistic audio samples. Following Karras et al. (2022) and the formulation in Moliner et al. (2024), we parameterize time by a continuous variable $\tau$, with $\mathbf{x}_{\tau}$ denoting the audio sample noisified at time $\tau$. At $\tau=T$, $\mathbf{x}_{T}$ is approximately Gaussian noise, and at $\tau=0$, $\mathbf{x}_{0}$ corresponds to clean audio. The probability-flow ODE for this process is given by
$d\mathbf{x}_{\tau}$ $=-\,\tau\,\nabla_{\mathbf{x}_{\tau}}\log\,p_{\tau}(\mathbf{x}_{\tau})\,d\tau,$ (20)
where $\nabla_{\mathbf{x}_{\tau}}\log\,p_{\tau}(\mathbf{x}_{\tau})$ is the prior score. We approximate this score using a denoiser network $s_{\theta}(\mathbf{x}_{\tau},\tau)$:
$\nabla_{\mathbf{x}_{\tau}}\log\,p_{\tau}(\mathbf{x}_{\tau})$ $\approx\frac{s_{\theta}(\mathbf{x}_{\tau},\tau)-\mathbf{x}_{\tau}}{\sigma(\tau)^{2}},$ (21)
where $\sigma(\tau)^{2}$ is the noise variance at time $\tau$. The network is trained via denoising score matching:
$\mathbb{E}_{\mathbf{x}_{0}\sim p_{\text{data}},\,\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})}\Bigl[\lambda(\tau)\,\bigl\|s_{\theta}(\mathbf{x}_{0}+\tau\bm{\epsilon},\tau)-\mathbf{x}_{0}\bigr\|_{2}^{2}\Bigr],$ (22)
with $\lambda(\tau)$ chosen as in Karras et al. (2022).
D.2 Methods
As described in the main text, for inverse problems with observations $\mathbf{y}$, we replace the prior score in equation 20 with the posterior score $\nabla_{\mathbf{x}_{\tau}}\log\,p_{\tau}(\mathbf{x}_{\tau}\mid\mathbf{y})$. Using a Bayesian decomposition Chung et al. (2022), we split the posterior score at time $\tau$ as
$\nabla_{\mathbf{x}_{\tau}}\log\,p_{\tau}(\mathbf{x}_{\tau}\mid\mathbf{y})$ $=\nabla_{\mathbf{x}_{\tau}}\log\,p_{\tau}(\mathbf{x}_{\tau})+\nabla_{\mathbf{x}_{\tau}}\log\,p_{\tau}(\mathbf{y}\mid\mathbf{x}_{\tau}),$ (23)
where the first term is the prior score from equation 21 and the second term is a likelihood term. In prior work Moliner et al. (2024), the likelihood score is approximated as the gradient of an audio-domain cost function $C_{\text{audio}}(\mathbf{y},H(\hat{\mathbf{x}}_{0}))$, scaled by a normalization factor $\xi(\tau)$:
$\nabla_{\mathbf{x}_{\tau}}\log p_{\tau}(\mathbf{y}\mid\mathbf{x}_{\tau})$ $\approx-\,\xi(\tau)\,\nabla_{\mathbf{x}_{\tau}}C_{\text{audio}}\bigl(\mathbf{y},\,H(\hat{\mathbf{x}}_{0})\bigr),\qquad\hat{\mathbf{x}}_{0}=D_{\theta}(\mathbf{x}_{\tau},\tau),$ (24)
where $H(\cdot)$ denotes the (known or learned) degradation operator (e.g., a frequency-domain equalizer in BABE2). Substituting equation 23 and equation 24 into the ODE equation 20 yields the posterior-flow:
$d\mathbf{x}_{\tau}$ $=-\,\tau\Bigl[\underbrace{\nabla_{\mathbf{x}_{\tau}}\log p_{\tau}(\mathbf{x}_{\tau})}_{\text{prior}}+\underbrace{\nabla_{\mathbf{x}_{\tau}}\log p_{\tau}(\mathbf{y}\mid\mathbf{x}_{\tau})}_{\text{likelihood}}\Bigr]d\tau.$ (25)
Moliner et al. (2024) uses a carefully designed approximation of the $\mathcal{A}$ operator and optimizes this quantity while carrying out diffusion inference steps. We sidestep this process via our contrastive framework.
Contrastive guidance. To incorporate CoGuide, we add a contrastive guidance term that measures alignment between the reconstruction and the observed degraded audio in an embedding space. We use an encoder, $f_{\varphi}:\mathcal{X}\cup\mathcal{Y}\to\mathbb{R}^{d}$, to map a candidate clean signal and the corresponding observation into a shared latent space. At diffusion time $\tau$, we obtain a current reconstruction $\hat{\mathbf{x}}_{0}=s_{\theta}(\mathbf{x}_{\tau},\tau)$ and compute embeddings $f_{\varphi}(\hat{\mathbf{x}}_{0}),f_{\varphi}(\mathbf{y})$ . We define a contrastive score via the inner product $\langle f_{\varphi}(\hat{\mathbf{x}}_{0})\,,f_{\psi}(\mathbf{y})\rangle$, with $\ell_{2}$-normalized embeddings to compute cosine similarity.
During sampling, this score acts as a surrogate likelihood (like in CoGuide): candidates whose embeddings align with the observation are favored. Concretely, we replace the likelihood term with a contrastive gradient given by:
$\widehat{\nabla}_{\mathbf{x}_{\tau_{k}}}\log p_{\tau_{k}}(\mathbf{y}\mid\mathbf{x}_{\tau_{k}})$ $\approx\gamma(\tau_{k})\,\nabla_{\mathbf{x}_{\tau_{k}}}\langle f_{\varphi}(\hat{\mathbf{x}}_{0})\,,f_{\varphi}(\mathbf{y})\rangle,$ (26)
where $\gamma(\tau)$ controls the guidance strength. The resulting CoGuide update becomes
$\mathbf{x}_{\tau_{k-1}}$ $=\mathbf{x}_{\tau_{k}}-h_{k}\,\tau_{k}\Bigl[\widehat{\nabla}_{\mathbf{x}_{\tau_{k}}}\log p_{\tau_{k}}(\mathbf{x}_{\tau_{k}})+\gamma(\tau_{k})\,\nabla_{\mathbf{x}_{\tau_{k}}}\langle f_{\varphi}(\hat{\mathbf{x}}_{0})\,,f_{\varphi}(\mathbf{y})\rangle)\Bigr],$ (27)
where the gradient $\nabla_{\mathbf{x}_{\tau_{k}}}$ is obtained by backpropagating the inner product $\langle f_{\varphi}(s_{\theta}(\mathbf{x}_{\tau_{k}},\tau_{k})),f_{\varphi}(\mathbf{y})\rangle$ through both the denoiser $s_{\theta}$ and the encoder $f_{\varphi}$.
D.3 Training and Implementation Details
Contrastive audio model. We train a contrastive encoder to embed clean and degraded piano audio into a shared latent space. Given a clean waveform $\mathbf{x}$ and its degraded counterpart $\mathbf{y}$, both sampled at 22,05022{,}050 Hz and cropped to fixed $8.35$ s segments, we first compute their complex Constant-Q Transform (CQT) Velasco et al. (2011).
$\mathbf{X}$ $=\mathrm{CQT}(\mathbf{x})\in\mathbb{C}^{F\times T},$ (28)
$\mathbf{Y}$ $=\mathrm{CQT}(\mathbf{y})\in\mathbb{C}^{F\times T},$ (29)
using $7$ octaves and $64$ bins per octave. Real and imaginary parts are stacked as channels and passed to the encoder $f_{\varphi}$. This encoder consists of a 2D CNN/ResNet backbone over the $(F,T)$ CQT grid, followed by global average pooling and a linear projection to a $d=256$-dimensional embedding, with $\ell_{2}$-normalization. Training. Clean data are drawn from the MAESTRO dataset Hawthorne et al. (2018), and degradations are applied on-the-fly in the waveform domain (e.g., low-pass filtering, spectral coloration, clipping, additive Gaussian noise, and mild dynamic range compression), producing pairs $(\mathbf{x}_{i},\mathbf{y}_{i})$. For a mini-batch, we compute similarities by inner products scaled by a temperature $\tau$. We use a symmetric InfoNCE loss so that each clean embedding treats its own degraded counterpart as the positive and all others as negatives (and vice versa). For a batch of size $B$, the loss is
$\mathcal{L}$ $=\frac{1}{2B}\sum_{i=1}^{B}\Biggl[-\log\frac{\exp\bigl(\langle f_{\varphi}(\mathbf{X}_{i}),f_{\varphi}(\mathbf{Y}_{i})\rangle/\tau\bigr)}{\sum_{j=1}^{B}\exp\bigl(\langle f_{\varphi}(\mathbf{X}_{i}),f_{\varphi}(\mathbf{Y}_{j})\rangle/\tau\bigr)}-\log\frac{\exp\bigl(\langle f_{\varphi}(\mathbf{Y}_{i}),f_{\varphi}(\mathbf{X}_{i})\rangle/\tau\bigr)}{\sum_{j=1}^{B}\exp\bigl(\langle f_{\varphi}(\mathbf{Y}_{i}),f_{\varphi}(\mathbf{X}_{j})\rangle/\tau\bigr)}\Biggr]$
The model is trained with AdamW (learning rate $10^{-4}$, weight decay 0), batch size $B=256$, and a fixed temperature $\tau=0.07$, until the validation contrastive loss and embedding alignment metrics saturate. The trained encoder $f_{\varphi}$ is then frozen and used to provide contrastive guidance during diffusion inference.
D.4 Results
Fig. 11 shows qualitative audio samples with spectrograms below for comparison across different composers’ pieces. CoGuide is able to reduce nonlinear effects, perform denoising of nonstationary noise, and enable the reconstruction of high-frequency details.
Figure 11: Comparison of spectrograms for the degraded (original), LTAS, and CoGuide across three different pieces of music from the MAESTRO dataset Hawthorne et al. (2018).1
Table 4 presents the quantitative metrics using Fréchet Audio Distance (FAD) Kilgour et al. (2018) (computed with VGGNet and PANN Kong et al. (2020a)) for the degraded recordings, the LTAS baseline, and CoGuide. Our method outperforms the baseline in both metrics. The LTAS baseline Moliner et al. (2024) is an equalization technique that seeks to align the Long-term Averaged Spectrum (LTAS) of the audio with a reference set. Additional qualitative audio samples are available at our project page.
Audio Type VGG $\downarrow$ PANN $\downarrow$
Degraded (Original Recording) 2.52 0.39
LTAS Baseline 2.88 0.27
CoGuide restored (Ours) 0.84 0.19
Table 4: Quantitative FAD comparison of CoGuide against the original recording and LTAS baseline for blind audio enhancement. FAD-PANN results are reported after scaling by $10^{3}$.
Appendix E Path-planning algorithms